# Robust Inference for Non-Linear Regression Models with Applications in Enzyme Kinetics

### v1

@@ 6.3 Tyrosinase enzyme data
7 Concluding remarks
In this paper we have proposed a robust method for parameter estimation under an NLR model based on the concept of minimum DPD estimation, introduced by Basu et al. [6]. We have established the desired asymptotic properties of our estimator and provided some mild conditions on the covariates and the regression function, under which those asymptotic properties hold. Their robustness has been verified by showing that these estimators have bounded influence functions. Under the Michaelis-Menten model in enzyme kinetics, extensive simulation study has been performed to empirically justify the robust performance of our method, which is then applied to robustly analyze several real-life datasets. Another fundamental paradigm of statistical inference, namely the hypothesis testing problem, has also been discussed in this paper. The MDPDE based Wald-type tests have been developed to test the significance of Michaelis-Menten model. The robustness of the testing procedure has also been validated through extensive simulation studies and theoretically via influence function analyses.
Our method can be treated as an optimization procedure with the minimum DPD objective function because, in most cases, minimizing the objective function is comparatively easier than solving the estimating equations. But this optimization is also a challenging problem because of the complexity of the objective function, particularly when the dimension of the regression parameter is high, where it may take a considerable amount of time. However, in our simulation experiments with the MM model, we have observed that the computation of the MDPDEs took less time than the existing robust methods. One also needs to carefully choose the initial parameter values to avoid the convergence of the optimization algorithm to a local minima. As a simple solution, one can run the algorithm for multiple initial values and finally take the solution with the lowest value of the objective function.
Finally, we would like to note that our NLR model is assumed to be homoscedastic and the errors are independently distributed. But these assumptions might not always be true in real-life situations involving NLR models. For example, some variants of the MM model (see, e.g., [33]) assume that the responses have constant coefficient of variation instead of having constant variance, implying heteroscedasticity in the assumed NLR model. So, an important future extension of our MDPDE based procedures developed in this paper would be to further extend them for such heteroscedastic NLR models as well as for the NLR models with dependent (e.g., autocorrelated) errors. We hope to pursue them in our sequel research works.
Appendix A Some background concepts
A.1 The DPD and the MDPDE
Basu et al. [6] proposed a robust method of estimation by minimizing the density power divergence (DPD) between the true data generating distribution (estimated suitably from the observed data) and the model distribution under the setup of IID data. Depending on a tuning parameter $\alpha\geq 0$, the DPD measure between two densities $f$ and $g$, defined with respect to a common dominating measure, is given by
$d_{\alpha}(g,f)=\left\{\begin{matrix}\int\left\{f^{1+\alpha}-\left(1+\frac{1}{\alpha}\right)f^{\alpha}g+\frac{1}{\alpha}g^{1+\alpha}\right\},&\mbox{if}&\alpha>0,\\ \int g\ln\left(\frac{g}{f}\right),&\mbox{if}&\alpha=0.\end{matrix}\right.$
Let $G$ be the (unknown) true data generating distribution with density $g$ and we want to model $g$ by the parametric family of densities =$\mathcal{F}=\left\{f_{\bm{\theta}}(\cdot):\bm{\theta}\in\bm{\Theta}\subseteq\mathbb{R}^{p}\right\}$. Now, in the minimum DPD approach, the most appropriate model density to fit $g$ well is that member of $\mathcal{F}$ for which $d_{\alpha}(g,f_{\bm{\theta}})$ is minimum over $\bm{\theta}\in\bm{\Theta}$. So, the minimum DPD functional at true distribution $G$ is defined by [6, 8]
[…]
Now, if an IID sample $x_{1},x_{2},\ldots,x_{n}$ is available from $G$, then the MDPDE $\widehat{\bm{\theta}}_{\alpha}$ of $\bm{\theta}$ can be obtained by replacing $G$ with the empirical distribution function $G_{n}$ in the above definition, leading to
$\widehat{\bm{\theta}}_{\alpha}=\bm{T}_{\alpha}(G_{n})=\arg\min_{\bm{\theta}\in\bm{\Theta}}\left\{\int f_{\bm{\theta}}^{1+\alpha}-\left(1+\frac{1}{\alpha}\right)\frac{1}{n}\sum_{i=1}^{n}f_{\bm{\theta}}^{\alpha}(x_{i})\right\}.$
Ghosh and Basu [14] generalized the definition of the MDPDE for independent non-homogeneous (INH) observations. Suppose we have $n$ independent observations $y_{1},y_{2},\ldots,y_{n}$ with $y_{i}\sim g_{i},\mbox{ for each }i=1,2,\ldots,n$, where $g_{i}$’s are possibly different densities and we wish to model $g_{i}$ by the parametric model family $\mathcal{F}_{i}=\left\{f_{i}(\cdot;\bm{\theta}):\bm{\theta}\in\bm{\Theta}\right\}$, for each $i=1,2,\ldots,n$. Note that, in this INH setup, although the distributions may be different, all of them must depend on the same parameter $\bm{\theta}$. Ghosh and Basu [14] defined the MDPDE of $\bm{\theta}$ as the minimizer of the average DPD measure between the data points and the corresponding models, given by
$H_{n}(\bm{\theta})=\frac{1}{n}\sum_{i=1}^{n}V_{i}(y_{i};\bm{\theta}),$ (36)
[…]
where $\bm{u}_{i}(y;\bm{\theta})=\nabla\ln f_{i}(y;\bm{\theta})$ denotes the score function.
Note that, as $\alpha\to 0$, the MDPDE is to be obtained by minimizing $\sum_{i=1}^{n}\left[-\ln\left(f_{i}(y_{i};\bm{\theta})\right)\right]$, or equivalently maximizing $\sum_{i=1}^{n}\ln\left(f_{i}(y_{i};\bm{\theta})\right)$ with respect to $\bm{\theta}$, or by solving the estimating equation $\sum_{i=1}^{n}\bm{u}_{i}(y_{i};\bm{\theta})=0$. So, the MDPDE coincides with the MLE at $\alpha=0$.
Considering the contaminated density for $i$-th observation being $g_{i,\epsilon}=(1-\epsilon)g_{i}+\epsilon\wedge_{t_{i}}$, where $\wedge_{t_{i}}$ is the density of degenerate distribution at the contamination point $t_{i}$ for $i=1,2,\ldots,n$, Ghosh and Basu [14] obtained the influence function of the MDPDE functional $\bm{T}_{\alpha}$. In particular, for contamination only in the $i_{0}$-th observation, this influence function turns out to be
$IF_{i_{0}}(t_{i_{0}},\bm{T}_{\alpha},\bm{G})=\bm{\Psi}_{n}^{-1}\frac{1}{n}\left[f_{i_{0}}(t_{i_{0}};\bm{\theta}^{g})^{\alpha}\bm{u}_{i_{0}}(t_{i_{0}};\bm{\theta}^{g})-\bm{\xi}_{i_{0}}\right],$ (38)
$IF(\bm{t},\bm{T}_{\alpha},\bm{G})=\bm{\Psi}_{n}^{-1}\frac{1}{n}\sum_{i=1}^{n}\left[f_{i}(t_{i};\bm{\theta}^{g})^{\alpha}\bm{u}_{i}(t_{i};\bm{\theta}^{g})-\bm{\xi}_{i}\right],$ (39)
[…]
A.2 Existing robust estimators under the NLR models
M-estimators:
For the NLR model given in (3), an M-estimator $\widehat{\bm{\beta}}_{M}$ of $\bm{\beta}$ is defined as
where $r_{i}=y_{i}-\mu(\bm{x_{i}},\bm{\beta}),\mbox{ for }i=1,\ldots,n$, $\bm{\Theta_{\beta}}$ is the parameter space of the regression parameter $\bm{\beta}$ and, if $\sigma$ is unknown, it needs to be estimated by any usual robust scale estimate, e.g., $\widehat{\sigma}=1.4826\mbox{Med}_{i}(|r_{i}-\mbox{Med}_{j}r_{j}|)$, and $\rho:\mathbb{R}\to\mathbb{R}^{+}$ is an appropriate even function which is non-decreasing in $\mathbb{R}^{+}$. When $\rho(\cdot)$ is differentiable and $\psi=\rho^{{}^{\prime}}$ then the associated estimating equation is given by
Following the asymptotic normality results of regression M estimators from Maronna et al. [29], it can be shown that under certain conditions the M-estimator $\bm{\widehat{\beta}}_{M}$, obtained by solving (41), satisfies
[…]
where $v_{M}=\sigma^{2}\frac{E\psi^{2}(\frac{u}{\sigma})}{\left[E\psi^{{}^{\prime}}(\frac{u}{\sigma})\right]^{2}},\mbox{ with }u\sim\mathcal{N}(0,\sigma^{2})$. This asymptotic variance formula has been used to compute the ARE for the M-estimator of $\bm{\beta}$, presented in Section 2.2.
The KPS estimator:
Tabatabai et al. [36] proposed a robust estimator of $\bm{\beta}$ under the NLR model in (3), by suitably modifying the M-estimation approach, as given by
where $\rho_{\omega}(x)=1-Sech(\omega x)$ with $Sech(x)=\frac{2}{e^{x}+e^{-x}}$, $t_{i}=\frac{1}{\sigma}(1-h_{ii})(y_{i}-\mu(\bm{x_{i}},\bm{\beta}))$, with $h_{ii}$ being the $i$-th diagonal element of the matrix $\bm{H}=\bm{\dot{\mu}(\dot{\mu}^{T}\dot{\mu})^{-1}\dot{\mu}^{T}}$ and $L_{i}=\sum_{j=1}^{k}\max\{M_{j},|x_{ij}|\},\ i=1,\ldots,n$, where $M_{j}=\mbox{Median}\{|x_{1j}|,\ldots,|x_{nj}|\},\ j=1,\ldots,k$, with $x_{ij}$ be the $j$-th element of the $k$-dimensional covariate vector $\bm{x_{i}}$ of the $i$-th observation.
If $\sigma$ is unknown, they suggested to use either of the two robust estimates defined as $\widehat{\sigma}=1.1926\mbox{ Med}_{1\leq i\leq n}\left(\mbox{Med}_{1\leq j\leq n}|r_{i}-r_{j}|\right)$ or $\widehat{\sigma}=2.2219\ \left\{|r_{i}-r_{j}|:1\leq i<j\leq n\right\}_{(l)}$, with $r_{i}=y_{i}-\mu\left(\bm{x_{i}},\bm{\widehat{\beta}}\right)$, $l=\dbinom{\left[\frac{n}{2}\right]+1}{2}$, the binomial coefficient, and $S_{(l)}$ denotes the $l$-th order statistic in the set $S$.
The MOM-estimator:
Liu et al. [26] proposed a very simple robust estimation method for the NLR model (3), called the median of means (MOM) method. This method consists of the following steps:
  • Obtain the OLS estimate of $\bm{\beta}$ from every group of observations separately, which we denote as $\widehat{\bm{\beta}}^{(j)}=\left(\widehat{\beta}_{1}^{(j)},\ldots,\widehat{\beta}_{p}^{(j)}\right)^{T}$ for group $j,\ j=1,2,\ldots,g$.
  • The MOM estimator is then given by $\widehat{\bm{\beta}}^{MOM}=\left(\widehat{\beta}_{1}^{MOM},\ldots,\widehat{\beta}_{p}^{MOM}\right)^{T}$, where $\widehat{\beta}_{k}^{MOM}=\mbox{Median}\left\{\widehat{\beta}_{k}^{(1)},\ldots,\widehat{\beta}_{k}^{(g)}\right\},\mbox{ for all }k=1,2,\ldots,p$.
Liu et al. [26] have also proved asymptotic properties (consistency and asymptotic normality) of the estimator under certain assumptions. However, from the definition of the method, it is intuitively clear that this method gives good, stable performance only when the sample size is moderately large and the proportion of outlier is small, which is also justified through our simulation studies, presented in the Section 4.3.
Appendix B Proof of theorems
B.1 Proof of Theorem 2.1
Lemma B.1.
Consider the setup of the NLR model (3) and suppose that the true data generating distributions belong to the assumed parametric model families. Then Conditions (R1) and (R2) imply Assumptions (A1)-(A7) of Ghosh and Basu [14].
Proof.
As the true distributions belong to the corresponding parametric model families, under which the distribution of $y_{i}$ is $\mathcal{N}\left(\mu(\bm{x_{i}},\bm{\beta}),\sigma^{2}\right)$ with $\mu(\bm{x_{i}},\bm{\beta})$ being thrice continuously differentiable, Assumptions (A1)-(A3) of Ghosh and Basu [14] directly follow from the property of the normal density in view of (R1). Assumption (A4) follows from Equation (14) of Condition (R2). The form of $V_{i}(y_{i};\bm{\theta})$ in (4) ensures that Assumption (A5) follows from Equation (13) of (R1). Now, to prove Equation (3.8) of Assumption (A6) of Ghosh and Basu [14], we note that
$\nabla_{j}V_{i}(y_{i};\bm{\theta})=\kappa e^{-\frac{\alpha}{2\sigma^{2}}(y_{i}-\mu(\bm{x_{i}},\bm{\beta}))^{2}}(y_{i}-\mu(\bm{x_{i}},\bm{\beta}))(\bm{\dot{\mu}})_{ij},~{}~{}~{}~{}j=1,\ldots,p,$
$=|\kappa|\frac{1}{n}\sum_{i=1}^{n}E_{i}\left[e^{-\frac{\alpha z_{i}^{2}}{2\sigma^{2}}}|z_{i}||(\bm{\dot{\mu}})_{ij}|\times I\left(e^{-\frac{\alpha z_{i}^{2}}{2\sigma^{2}}}|z_{i}||(\bm{\dot{\mu}})_{ij}|>\frac{N}{|\kappa|}\right)\right],\mbox{ letting }z_{i}=y_{i}-\mu(\bm{x_{i}},\bm{\beta})$
$\leq|\kappa|\frac{1}{n}\sum_{i=1}^{n}|(\bm{\dot{\mu}})_{ij}|E_{i}\left[e^{-\frac{\alpha z_{i}^{2}}{2\sigma^{2}}}|z_{i}|\times I\left(e^{-\frac{\alpha z_{i}^{2}}{2\sigma^{2}}}|z_{i}|>\frac{N}{|\kappa|(\sup_{n>1}\max_{1\leq i\leq n}|(\bm{\dot{\mu}})_{ij}|)}\right)\right]$
$=|\kappa|E_{1}\left[e^{-\frac{\alpha z_{1}^{2}}{2\sigma^{2}}}|z_{1}|\times I\left(e^{-\frac{\alpha z_{1}^{2}}{2\sigma^{2}}}|z_{1}|>\frac{N}{|\kappa|(\sup_{n>1}\max_{1\leq i\leq n}|(\bm{\dot{\mu}})_{ij}|)}\right)\right]\left(\frac{1}{n}\sum_{i=1}^{n}|(\bm{\dot{\mu}})_{ij}|\right),$
as $z_{i}$’s are IID. Since we have $\sup_{n>1}\max_{1\leq i\leq n}|(\bm{\dot{\mu}})_{ij}|=O(1)$ by (12) of (R1), so, by dominated convergence theorem (DCT)
$\lim_{N\to\infty}E_{1}\left[e^{-\frac{\alpha z_{1}^{2}}{2\sigma^{2}}}|z_{1}|\times I\left(e^{-\frac{\alpha z_{1}^{2}}{2\sigma^{2}}}|z_{1}|>\frac{N}{|\kappa|(\sup_{n>1}\max_{1\leq i\leq n}|(\bm{\dot{\mu}})_{ij}|)}\right)\right]=0.$
We also have
$\sup_{n>1}\left(\frac{1}{n}\sum_{i=1}^{n}|(\bm{\dot{\mu}})_{ij}|\right)\leq\sup_{n>1}\max_{1\leq i\leq n}|(\bm{\dot{\mu}})_{ij}|=O(1).$
Hence, Equation (3.8) of (A6) follows for all $j=1,2,\ldots,p$. Similarly one can show that it follows also for $j=p+1$, i.e., for $\nabla_{\sigma^{2}}$, the gradient with respect to $\sigma^{2}$. Similarly, Equation (3.9) of Assumption (A6) and Assumption (A7) of Ghosh and Basu [14] can also be shown to hold using second part of (12) of (R1) and (15) of (R2), respectively. ∎
Proof of the theorem:
The consistency part follows from Lemma B.1 and Theorem 3.1 of Ghosh and Basu [14]. Also, using Lemma (B.1), it follows from Theorem 3.1 of Ghosh and Basu [14] that the asymptotic distribution of $\bm{\Omega}_{n}^{-\frac{1}{2}}\bm{\Psi}_{n}\left[\sqrt{n}(\bm{\widehat{\theta}_{\alpha}}-\bm{\theta}_{0})\right]$ is $\mathcal{N}_{p+1}(\bm{0},\mathbb{I}_{p+1})$, where $\bm{\Omega}_{n}$ and $\bm{\Psi}_{n}$ are as defined in Equations (3.3) and (3.4) in Ghosh and Basu [14], and are to be computed at the model density with $\bm{\theta}=\bm{\theta_{0}}$. For the present case of NLR model (3), we can see that
$\bm{u}_{i}(y;\bm{\theta})=\nabla_{\bm{\theta}}\ln f_{i}(y;\bm{\theta})=\begin{pmatrix}\frac{(y-\mu(\bm{x_{i}},\bm{\beta}))}{\sigma^{2}}\nabla_{\bm{\beta}}\mu(\bm{x_{i}},\bm{\beta})\\ \\ \frac{(y-\mu(\bm{x_{i}},\bm{\beta}))^{2}}{2\sigma^{4}}-\frac{1}{2\sigma^{2}}\end{pmatrix}_{(p+1)\times 1}.$ (42)
[…]
So, a simplified form of the matrix $\bm{\Psi}_{n}$ is given by
$\bm{\Psi}_{n}=\frac{1}{n}\sum_{i=1}^{n}\bm{J}_{\alpha}^{(i)}=\begin{bmatrix}\frac{\zeta_{\alpha}}{n}\left(\bm{\dot{\mu}(\beta)}^{T}\bm{\dot{\mu}(\beta)}\right)&\bm{0}_{p}\\ \\ \bm{0}_{p}^{T}&\varsigma_{\alpha}\end{bmatrix}.$
We also have
[…]
and using it we obtain a simplified form of the matrix $\bm{\Omega}_{n}$ as
$\bm{\Omega}_{n}=\frac{1}{n}\sum_{i=1}^{n}(\bm{J}^{(i)}_{2\alpha}-\bm{\xi}_{i}\bm{\xi}_{i}^{T})=\begin{bmatrix}\frac{\zeta_{2\alpha}}{n}\left(\bm{\dot{\mu}(\beta)}^{T}\bm{\dot{\mu}(\beta)}\right)&\bm{0}_{p}\\ \\ \bm{0}_{p}^{T}&\varsigma_{2\alpha}-\frac{\alpha^{2}}{4}\zeta_{\alpha}^{2}\end{bmatrix}.$
$\bm{\Omega}_{n}^{-\frac{1}{2}}\bm{\Psi}_{n}\left[\sqrt{n}\left(\bm{\widehat{\theta}_{\alpha}}-\bm{\theta_{0}}\right)\right]=\begin{pmatrix}\frac{\zeta_{\alpha}}{\sqrt{\zeta_{2\alpha}}}\left(\bm{\dot{\mu}(\beta_{0})}^{T}\bm{\dot{\mu}(\beta_{0})}\right)^{1/2}\left(\bm{\widehat{\beta}_{\alpha}}-\bm{\beta_{0}}\right)\\ \\ \frac{\varsigma_{\alpha}}{\sqrt{\varsigma_{2\alpha}-\frac{\alpha^{2}}{4}\zeta_{\alpha}^{2}}}\sqrt{n}\left(\widehat{\sigma}_{\alpha}^{2}-\sigma_{0}^{2}\right)\end{pmatrix},$
which asymptotically follows $\mathcal{N}_{p+1}(\bm{0},\mathbb{I}_{p+1})$, where $\zeta_{p}$’s and $\varsigma_{p}$’s are to be computed at $\sigma=\sigma_{0}$. Therefore, asymptotically, $\frac{\zeta_{\alpha}}{\sqrt{\zeta_{2\alpha}}}\left(\bm{\dot{\mu}(\beta_{0})}^{T}\bm{\dot{\mu}(\beta_{0})}\right)^{1/2}\left(\bm{\widehat{\beta}_{\alpha}}-\bm{\beta_{0}}\right)$ and
$\frac{\varsigma_{\alpha}}{\sqrt{\varsigma_{2\alpha}-\frac{\alpha^{2}}{4}\zeta_{\alpha}^{2}}}\sqrt{n}\left(\widehat{\sigma}^{2}-\sigma_{0}^{2}\right)$ are independently distributed with their respective asymptotic distributions being $\mathcal{N}_{p}(\bm{0,\mathbb{I}_{p}})$ and $\mathcal{N}(0,1)$. This completes the proof.
B.2 Proof of Proposition 1
For the case of the MM model in (28), $\bm{\dot{\mu}(\beta)}$ (given in (29)) is of order $n\times 2$ and
$\sup_{n>1}\max_{1\leq i\leq n}|(\bm{\dot{\mu}})_{i1}|=\sup_{n>1}\max_{1\leq i\leq n}\frac{x_{i}}{\beta_{2}+x_{i}},$ (45)
$\sup_{n>1}\max_{1\leq i\leq n}|(\bm{\dot{\mu}})_{i2}|=|\beta_{1}|\sup_{n>1}\max_{1\leq i\leq n}\frac{x_{i}}{(\beta_{2}+x_{i})^{2}},$ (46)
$\sup_{n>1}\max_{1\leq i\leq n}|(\bm{\dot{\mu}})_{i1}(\bm{\dot{\mu}})_{i2}|=|\beta_{1}|\sup_{n>1}\max_{1\leq i\leq n}\frac{x_{i}^{2}}{(\beta_{2}+x_{i})^{3}}.$ (47)
We have $x_{i}>0$ for all $i$ and $\beta_{2}\geq 0$. So, $\frac{x_{i}}{\beta_{2}+x_{i}}$ is bounded between 00 and $1$. So, the quantity in Equation (45) is $O(1)$.
and
B.3 Proof of Proposition 2
Let us define $A_{n}=\frac{1}{n}\bm{\dot{\mu}(\beta)}^{T}\bm{\dot{\mu}(\beta)}$. The characteristic equation of $A_{n}$ is
$\lambda^{2}-tr(A_{n})\lambda+det(A_{n})=0.$
So, the eigenvalues of $A_{n}$ are given by
$\lambda=\frac{tr(A_{n})\pm\sqrt{tr(A_{n})^{2}-4det(A_{n})}}{2}.$
Here $\lambda$’s are real as $A_{n}$ is real symmetric matrix. Now, in order to have the minimum eigenvalue of $A_{n}$ strictly positive, we must have
$tr(A_{n})-\sqrt{tr(A_{n})^{2}-4det(A_{n})}>0$
$\Leftrightarrow det(A_{n})>0$
$\Leftrightarrow\sum_{i=1}^{n}a_{i}^{2}\ \ \sum_{i=1}^{n}b_{i}^{2}>\left(\sum_{i=1}^{n}a_{i}b_{i}\right)^{2},$
with $a_{i}=\frac{x_{i}}{\beta_{2}+x_{i}},b_{i}=\frac{\beta_{1}x_{i}}{(\beta_{2}+x_{i})^{2}}$, which is true by Cauchy-Schwartz inequality unless the equality occurs. The equality occurs only when $a_{i}\propto b_{i}\ \mbox{ for all }i$, i.e., when $x_{i}=$ constant, for all $i$ or $\beta_{1}=0$.
Therefore, in order to have the minimum eigenvalue of $A_{n}$ strictly positive and satisfy (14), it is necessary that the values $x_{i}$’s are not all same, for $n\geq 2$, and $\beta_{1}\neq 0$.
Appendix C Additional numerical results
C.1 Influence functions
[…]
(a) IF of $T_{\alpha}^{\beta_{1}}$ with $i_{0}=10$
(b) IF of $T_{\alpha}^{\beta_{2}}$ with $i_{0}=10$
(c) IF of $T_{\alpha}^{\beta_{1}}$ with $i_{0}=40$
(d) IF of $T_{\alpha}^{\beta_{2}}$ with $i_{0}=40$
Figure 7: IFs of the MDPDEs of the parameters $\beta_{1}$ and $\beta_{2}$ in the MM model, for contamination only in $i_{0}$-th observation under Setup 1
(a) IF of $T_{\alpha}^{\beta_{1}}$ with $i_{0}=15$
(b) IF of $T_{\alpha}^{\beta_{2}}$ with $i_{0}=15$
(c) IF of $T_{\alpha}^{\beta_{1}}$ with $i_{0}=45$
(d) IF of $T_{\alpha}^{\beta_{2}}$ with $i_{0}=45$
Figure 8: IFs of the MDPDEs of the parameters $\beta_{1}$ and $\beta_{2}$ in the MM model, for contamination only in $i_{0}$-th observation under Setup 2
C.2 Additional tables from simulation results
Table 10 presents the average prediction errors for the fits obtained by the MDPDE and its robust competitors under the MM model (28), based on 1000 replications for contaminated data with contamination proportions $e_{c}=20\%$ and $30\%$; associated EBias and EMSE of the parameter estimates are given in Tables 11 and 12, respectively.
Setup Setup 1 Setup 2
Outlier Prop 20% 30% 20% 30% 20% 30% 20% 30%
OLS 15.46 32.75 8.214 11.59 79.51 170.43 100.69 172.67
KPS(w=0.5) 1.094 13.02 0.976 2.009 3.802 3.997 3.810 3.901
KPS(w=1) 1.041 2.745 1.080 2.258 3.827 3.800 3.829 3.809
KPS(w=1.5) 1.208 1.642 1.010 1.214 3.889 3.838 3.891 3.843
Huber 1.221 2.408 2.784 9.485 5.008 11.15 14.18 127.95
Tukey 0.947 0.941 0.947 1.829 5.930 5.548 14.81 124.12
MDPDE($\alpha$=0.05) 12.93 30.06 9.498 13.37 68.41 158.58 88.67 165.00
MDPDE($\alpha$=0.1) 10.37 27.32 8.156 11.74 54.83 162.34 80.34 181.05
MDPDE($\alpha$=0.2) 1.500 21.77 3.735 9.261 3.830 19.97 5.486 96.12
MDPDE($\alpha$=0.3) 0.953 4.785 0.950 5.183 7.194 4.564 3.800 8.863
MDPDE($\alpha$=0.5) 0.956 0.950 0.954 0.946 3.811 3.794 3.816 3.807
MDPDE($\alpha$=0.7) 0.962 0.951 0.961 0.949 3.829 3.808 3.834 3.820
MDPDE($\alpha$=1) 0.972 0.956 0.971 0.955 3.856 3.826 3.862 3.838
Outlier proportion 10% 40% 10% 40%
Empirical Bias
OLS 4.9733 2.6438 7.2248 2.7391 56.40 260.74 42.21 163.89
MOM 4.1274 3.9194 124.75 7.3124 3.1807 221.13 5.1049 356.55
KPS(w=0.5) 0.0132 0.0094 0.0221 0.1983 0.0521 0.1183 1.1789 4.7107
KPS(w=1) 0.0234 0.1011 0.0115 0.0457 0.0287 0.1117 0.0314 0.0955
KPS(w=1.5) 0.0388 0.1583 0.0247 0.1038 0.0405 0.1605 0.0335 0.1268
Empirical MSE
OLS 36.43 99.85 83.95 192.01 7559 175837 1866 29046
MOM 49.79 612.19 6.8E+06 1737.08 12.67 136261 55.56 5.7E+06
KPS(w=0.5) 0.0712 0.5051 0.1184 0.8523 0.0700 0.4974 50.62 1037
KPS(w=1) 0.0757 0.5376 0.0864 0.6142 0.0740 0.5260 0.0822 0.5970
KPS(w=1.5) 0.0941 0.7241 0.0971 0.7193 0.0943 0.7115 0.0935 0.6786
M-Est (Huber) 0.3977 0.5981 2.4094 2.2897 10.40 79.48 34.58 81.00
Outlier proportion 10% 40% 10% 40%
Empirical Bias
OLS 10.54 0.2015 15.72 0.24382 28.55 7.6917 41.55 11.09
MOM 8.3100 0.4567 15.99 0.7324 28.70 6.6655 47.51 11.22
KPS(w=0.5) 0.0666 0.0289 0.2759 0.1323 0.0822 0.0411 0.1605 0.1090
KPS(w=1) 0.0003 0.0016 0.0134 0.0109 0.0180 0.0094 0.0575 0.0309
KPS(w=1.5) 0.0186 0.0042 0.0198 0.0014 0.0050 0.0011 0.0125 0.0117
M-Est (Huber) 0.8973 0.0919 2.2299 0.1811 4.8576 1.0918 27.05 5.8772
M-Est (Tukey) 0.0685 0.0031 0.2401 0.0128 3.5690 0.9716 26.05 5.7811
MDPDE($\alpha$=0.05) 0.5644 0.5019 5.3485 0.3564 1.2477 12.89 7.3678 1.3647 7.1999 16.44 3.5317 14.66
MDPDE($\alpha$=0.1) 0.0341 0.0288 0.0947 0.2141 0.0747 0.7250 0.0764 0.0528 0.1385 1.7809 0.2961 1.4378
MDPDE($\alpha$=0.2) 0.0119 0.0059 0.0054 0.0125 0.0108 0.0601 0.0422 0.0199 0.0225 0.0797 0.0419 0.0971
MDPDE($\alpha$=0.3) 0.0033 0.0022 0.0277 0.0051 0.0036 0.1037 0.0251 0.0120 0.0368 0.0419 0.0244 0.1232
MDPDE($\alpha$=0.5) 0.0014 0.0003 0.0840 0.0142 0.0000 0.2200 0.0131 0.0067 0.0873 0.0166 0.0132 0.2280
MDPDE($\alpha$=0.7) 0.0039 0.0007 0.1416 0.0191 0.0019 0.3398 0.0076 0.0043 0.1418 0.0077 0.0094 0.3419
MDPDE($\alpha$=1) 0.0071 0.0018 0.2163 0.0219 0.0029 0.4972 0.0029 0.0025 0.2136 0.0000 0.0063 0.4942
Empirical MSE
OLS 127.86 1.2657 268.46 1.3927 1026.58 104.58 1993.45 181.41
MOM 154.28 3.5655 383.09 6.8593 1324.69 119.51 2784.57 227.17
KPS(w=0.5) 0.4331 0.0375 0.8790 0.1089 0.4197 0.0382 0.5939 0.0709
KPS(w=1) 0.4821 0.0392 0.5435 0.0437 0.4766 0.0391 0.5443 0.0469
KPS(w=1.5) 0.6142 0.0497 0.6296 0.0495 0.6076 0.0495 0.6297 0.0518
M-Est (Huber) 1.7325 0.1060 12.83 0.4652 91.42 6.7782 912.82 45.14
M-Est (Tukey) 2.9843 0.0343 9.4442 0.2652 104.27 7.8639 892.73 44.69
MDPDE($\alpha$=0.05) 63.26 11.48 102.59 151.63 8.7679 365.28 294.77 49.97 116.81 918.56 83.04 370.88
MDPDE($\alpha$=0.1) 0.4434 0.0463 1.0514 5.3745 0.2249 12.67 0.7096 0.0522 0.8664 65.63 4.1487 22.13
MDPDE($\alpha$=0.2) 0.4017 0.0345 0.0593 0.4890 0.0428 0.0828 0.4047 0.0357 0.0641 0.5229 0.0522 0.1032
MDPDE($\alpha$=0.3) 0.4112 0.0350 0.0638 0.4857 0.0409 0.0909 0.4146 0.0358 0.0663 0.5135 0.0470 0.0996
MDPDE($\alpha$=0.5) 0.4451 0.0381 0.0828 0.5136 0.0432 0.1449 0.4473 0.0385 0.0839 0.5264 0.0460 0.1489
MDPDE($\alpha$=0.7) 0.4821 0.0415 0.1088 0.5412 0.0455 0.2289 0.4848 0.0419 0.1090 0.5536 0.0481 0.2300
MDPDE($\alpha$=1) 0.5374 0.0463 0.1521 0.5806 0.0489 0.3834 0.5420 0.0470 0.1517 0.5934 0.0512 0.3794

### v2

@@ D.1 Tyrosinase enzyme data
The IWJ method was used for the selection of the optimum tuning parameter value for these datasets. For Agaricus bisporus (Ab), it took 5 iterations to converge to the optimum value $\widehat{\alpha}=0.35$, while for Pleurotus ostreatus (Po) data the method required only one iteration to select the same optimum value $\widehat{\alpha}=0.53$ for all pilot choices of $\alpha$.
The scatter plot of Ab data is provided in the Figure 9(a) along with the MDPDE fits (for various $\alpha\geq 0$) and the robust competitors. The figure shows that the proposed MDPDE (for all $\alpha\geq 0$), as well as all robust competitors, provide good fit. However, Figure 9(b), corresponding to the Po dataset, shows that the MDPDEs with $\alpha\geq 0.53$ and KPS with $w=1$ provide good robust fits with significantly lower TAPE than other robust competitors.
We further tested the validity of the MM model for these data using the Wald-type tests, which resulted a p-value of <0.001<0.001 for all values of $\alpha\geq 0$ (including the classical Wald test). It shows that the assumed MM model is valid, and all tests give the same inference in the presence of no or mild contamination.
D.2 A putative enzyme data
This dataset is taken from Table 3.1 of Maragoni [30], containing the reaction velocities of an enzyme kinetic reaction at 15 different concentration levels of a putative enzyme, recorded over 5 different replications. So, there are 75 data points in total, making it a large sample in the context of enzyme kinetics. We have fitted the MM model to this dataset using the proposed MDPDE, as well as its existing competitors. The parameter estimates are given in Table 13 along with their asymptotic standard errors and $20\%$ TAPEs for the fitted models, while the model fits are illustrated in Figure 10. The optimum value of $\alpha$ in the MDPDEs, for this dataset, was observed to converge to $0.49$ (starting with the MDPDE with $\alpha>0.4$), using the IWJ method. One can clearly observe that the proposed MDPDE at optimum $\alpha=0.49$ (in fact, for all $\alpha\geq 0.49$) yields robust estimates and improved model fits, ignoring the effects of the outliers.
We also fitted separate MM models to the individual replications using all other robust methods under discussion. However, no significant variations in the results were observed across replications, and hence they are omitted here for brevity; only the results from the combined sample of size 75 are reported (as an illustration of the MDPDEs under larger sample sizes).
Table 13: Parameter estimates of the MM model fitted to the putative enzyme data, along with their asymptotic standard errors (in parentheses) and $20\%$ TAPE. [For MDPDE, $\widehat{\alpha}$ near $0.49$ is the optimum choice, obtained by IWJ method]
MDPDE
$\alpha$ 0 0.05 0.1 0.2 0.3 0.49 0.7 1
$\beta_{1}$ 81.098 80.569 80.053 79.145 78.575 78.126 77.858 77.313
(1.985) (2.622) (2.552) (2.419) (2.331) (2.298) (2.342) (2.428)
$\beta_{2}$ 38.617 38.350 38.098 37.693 37.525 37.529 37.506 37.092
(2.413) (3.193) (3.114) (2.964) (2.868) (2.845) (2.907) (3.013)
$\sigma$ 3.831 5.074 4.938 4.643 4.399 4.147 4.007 3.876
(2.397) (4.219) (4.032) (3.669) (3.417) (3.273) (3.291) (3.335)
TAPE 8.836 8.659 8.411 8.039 7.900 7.831 7.787 7.765
KPS M-est M-est
w = 0.5 w = 1 w = 1.5 Huber Tukey
$\beta_{1}$ 79.538 79.229 79.451 79.416 78.446
(2.118) (2.504) (3.100) (2.771) (2.186)
$\beta_{2}$ 38.270 38.488 38.736 38.040 37.320
(2.609) (3.108) (3.854) (3.405) (2.685)
TAPE 8.046 7.922 7.948 8.055 7.906
Figure 10: The putative enzyme data and the plot of the fitted MM models using different methods